OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5.

OpenStudy (anonymous):

y=-7x-5 m=-7 (y-y1)=m(x-x1) (y+3)=_7(x-1)

OpenStudy (anonymous):

(y+3)=-7(x-1) then you will get y=-7x+4

OpenStudy (kymber):

Since it's parallel to the line, they'll have the same slope plug in the numbers you know & solve for B